Could someone please take the time to help me solve the question: Solve the following system of two linear equations by elimination: 3x – 4y = 13 and
3y + 2 x = 3.
Thank you, I really appreciate your help.

Answer by ankor@dixie-net.com(22740)

Solve the following system of two linear equations by elimination:
3x – 4y = 13
3y + 2x = 3
:
Arrange the 2nd equation to be in the standard form like the 1st equation:
3x - 4y = 13
2x + 3y = 3
:
We will eliminate x,
multiply the 1st equation by 2,
multiply the 2nd equation by 3
:
6x - 8y = 26
6x + 9y = 9
---------------subtracting eliminates x, find y
0x - 17y = 17
:
-17y = 17
y = 17/-17; divide both sides by -17
y = -1
:
Substitute -1 for y in the 1st equation and find x
3x - 4(-1) = 13
3x + 4 = 13; minus a minus is a plus
3x = 13 - 4
3x = 9
x = 9/3
x = +3
:
:
Check both solutions by substitution in the 2nd equation:
3y + 2x = 3
3(-1) + 2(3) = 3
-3 + 6 = 3; confirms our solutions
:
Could you follow this OK?